OpenStudy (anonymous):

Which of the following represents x^2-16/x^2-4x in simplified form?

mathslover (mathslover):

did u mean this ?: \[\large{\frac{x^2-16}{x^2-4x}}\]

mathslover (mathslover):

@jrosesweet

OpenStudy (anonymous):

Yes :) @mathslover

mathslover (mathslover):

\[\large{\frac{(x+4)(x-4)}{x(x-4)}}\] right?

mathslover (mathslover):

since : \[\large{a^2-b^2=(a+b)(a-b)}\]

mathslover (mathslover):

now cancel out x-4 \[\large{\frac{(x+4)\cancel{(x-4)}^1}{x\cancel{(x-4)}^1}}\] \[\large{\frac{x+4}{x}}\]

mathslover (mathslover):

\[\large{\frac{x}{x}+\frac{4}{x}}\] \[\large{1+\frac{4}{x}}\]

mathslover (mathslover):

choices given ?@jrosesweet

OpenStudy (anonymous):

You got it right that was one of the choices!

mathslover (mathslover):

:D i am mathslover dude ..